OpenStudy (anonymous):

the perimeter of a rectangle is cm. one side is 11cm longer than other side. find the lenghts of the sides,

OpenStudy (ajprincess):

The perimeter is not given.

OpenStudy (anonymous):

34cm

Parth (parthkohli):

If one side is x cm, then the other will be x + 11 The formula for perimeter of rectangle is 2(l + b) => 2(x + x + 11) = 34 Solve for x.

OpenStudy (campbell_st):

let x = width then x + 11 = length P = 2 *length + 2 * width 24 = 2x + 2(x + 11) 24 = 4x + 22 solve the equation for x... the width then find the length to get the dimensions

OpenStudy (campbell_st):

oops 34 = 4x + 22

OpenStudy (ajprincess):

Say x and y as the length and width of the rectangle. x+y+x+y=perimeter. Since it is given that one side is 11cm longer than the other x=y+11 11+y+y+11+y+y=34 4y+22=34 4y=34-22 4y=12 y=3 x=11+3 x=14

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

yw